OpenStudy (jenny.andrade):

Which of the following is the correct application of the quadratic formula for the equation 6x^2-7x+2 = 0 ? Select one: a. x=(-6+-sqrt(36+56))/(4) b. x=(7+-sqrt(49-48))/(12) c. x=(-6+-sqrt(36+56))/(-7) d. x=(7+-sqrt(49+48))/(-12) i dont know what quadratic formula is?

OpenStudy (mathstudent55):

The solution to the equation \(ax^2 + bx + c = 0\) is \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)